OpenStudy (anonymous):

A bag contains 2 steel balls and 5 brass balls. The total weight is 13 pounds. If 2 steel balls are added and 2 brass balls are removed, the weight decreases to 12 pounds. How much does each kind of ball weigh?

OpenStudy (anonymous):

@Vandreigan can you help with this one?

OpenStudy (anonymous):

let x be steel balls let y be brass balls 2x+5y=13 4x+3y=12 solve simulteneously

OpenStudy (anonymous):

I don't know how to do that :/ I'm in a GED class, I never learned this.

OpenStudy (anonymous):

We will call our equations 2x+5y=13 i 4x+3y=12 ii we multiply equation i by 2 to get equation 3 4x+10y=26 iii

OpenStudy (anonymous):

We then subtract equation 3 from 2 4x+3y=12 ii - 4x+10y=26 iii -7y=-14

OpenStudy (anonymous):

y becomes 2 we sub 2 for y in equation 1 2x +(5*2)=13 and solve for x

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

1.5, correct?

OpenStudy (anonymous):

yep x=1.5

OpenStudy (anonymous):

2, yes?

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

m bad i realised we sloved earlier

OpenStudy (anonymous):

its okay!

OpenStudy (anonymous):

anytime